10. If the price of face mask is increasing from 5 birr to 25 birr the quantity demand for face maskis decreasing from 100 to 20 birr. Calculate price elasticity of demand?
1
Expert's answer
2020-04-02T09:52:11-0400

Quantity mid point = "(Q1 + Q2)\/2" = (100 + 20)/2 = 60

Price mid point = "(P1 + P2)\/2" = (5 + 25)/2 = 15

% change in quantity = "(Q2 - Q1)\/Quantity mid point" = (20 - 100)/60 = -1.33

% change in price = "(P2 - P1)\/Price mid point" = (25 - 5)/15 = 1.33

Arc elasticity of demand = -1.33/1.33 = 1